dwvdsv:

A rectangle is placed around a semicircle as shown below. The width of the rectangle is 2 yd. Find the area of the shaded region. (shaded region is the rectangle) Use the value 3.14 for π , and do not round your answer. Be sure to include the correct unit in your answer.

dwvdsv:

Florisalreadytaken:

we first should find \( \frac{1}{2} \) of the area of the WHOLE circle, after that find the area of the rectangle, and then subtract them

1 attachment
Florisalreadytaken:

lets find the area of the circle first \[ Area_\color{lightskyblue}{o} = \pi r^2 \rightarrow Area_\color{lightskyblue}{o} =3.14 \times 2^2 \rightarrow Area_\color{lightskyblue}{o} = 12.56 \]

Florisalreadytaken:

can you do the rectangle?